# Itô’s formula for flows of conditional measures on semimartingales

### v2

@@ 5.1. Construction of the conditional independent copy
where the fourth equation holds since ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equation holds because of the same calculation as (5.5) and the last equation holds because of (2) in Proposition 5.1. Therefore,
¯$\overline{\mathbb{P}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{\mathbb{P}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{\mathbb{P}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}),$
and thus completes the proof.∎
5.2. Proof of Theorem 3.3
For (1), due to the decomposablity (see [37, Theorem 1, Chapter 4]) let us assume that
@@
where $M_{t}$ is an ${\mathbb{F}}$-local martingale and $V_{t}$ is finite variation process. To prove that $X$ is a $\overline{\mathbb{F}}$-semimartingale, it suffices to show that $M_{t}$ is a $\overline{\mathbb{F}}$-local martingale. Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of ${\mathbb{F}}$-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an ${\mathbb{F}}$-martingale for all $k\in{\mathbb{N}}^{+}$ and as $k\to\infty$, we have $\tau_{k}\to\infty$. We claim that
@@
¯$\overline{\mathcal{F}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
@@
for all $A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}$. Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditional independent of given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
@@
By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$, we know that
@@
$=\overline{\mathbb{E}}\Big{[}\overline{\mathbb{E}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{\mathbb{P}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{\mathbb{P}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]}$
where the second inequality holds by the tower rule. Since $M^{\tau_{k}}$ is a ${\mathcal{F}}$-martingale, we get
@@
Repeating the calculation as in (5.7), (5.8) and (5.9) we achieve
@@
To show that $X^{\prime}$ is an $\overline{\mathbb{F}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditional independent copy of $M,V$, and $\tau_{k}$. Then it suffices to show that it suffices to show that
@@
Similarly as the procedure of showing (5.6), we have
@@
where the first and last equation mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equation hold since (2) in Proposition 5.1, the third and the fifth equation is because of the tower rule and the fourth equation follows from the martingale property. This completes the proof of (1).
@@
For (2b), due to (2a) and the decomposablity (see [37, Theorem 1, Chapter 4]), it suffices to show that if $X$ is a ${\mathbb{F}}$-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{\mathbb{F}}$, ${\mathbb{F}}$ and ${\mathbb{G}}$ and thanks to (1), it can be further reduced to prove that $X^{\mathcal{G}}$ is a local martingale with respect to ${\mathbb{F}}$ and ${\mathbb{G}}$. Let $\tau_{k}$ be the stopping time $\tau_{k}:=\inf\{t\in[0,T]:[X,X]_{t}>k\}$ and consider the truncated process $X^{\tau_{k}}:=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$. By Burkholder-Davis-Gundy inequality,
@@
for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the ${\mathbb{F}}$-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$. Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ are ${\mathbb{F}}$-martingales for all $\ell\in{\mathbb{N}}^{+}$. Using conditional dominate convergence theorem, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$ and (5.10), we have
@@
Combining with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable ${\mathbb{F}}$-martingale.
Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a ${\mathbb{G}}$-martingale, and using the property of H hyphothesis, and the compatiability condition of ${\mathbb{F}}$ and ${\mathbb{G}}$ (i.e. ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$ for all $t\in[0,T]$), it is also an ${\mathbb{F}}$-martingale (see [20]). The integrability holds trivially since $X_{t}$ is integrable. And the martingale property holds since
@@
This completes the proof of the claim that $X^{\mathcal{G}}$ is an ${\mathbb{G}}$-martingale and thus leads to(2b).
For (2c), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity (the definition of a sequence tending to identity can be found in the [37, Page 64, defintion] with a slight modification of the terminal time), such that
@@
where the convergence is understood as almost surely convergence. Such partition can be found thanks to [37, Theorem 21, Chapter II]. Then
@@
Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$ where $M^{\prime}$ is a local martingale and $V^{\prime}$ being a adapted finite variation process and
$\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$
Therefore,
@@
For the first term in (5.11), we use Burkholder-Davis-Gundy inequality so that
@@
$\overline{\mathbb{E}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{\mathbb{E}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{\mathbb{E}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty$
Since $Z$ is a càdlàg function, we know that $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$, and we can use the dominated convergence theorem (the classical one and the one in ${\mathcal{H}}_{p}$, see [37, Page 273, Lemma]) to get
@@
Noting that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
@@
¯$\overline{\mathbb{P}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{\mathbb{E}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
@@
¯$\overline{\mathbb{P}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
Therefore $R_{m}\to R$ in probability. Meanwhile $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ almost surely. Therefore,
@@
For (2d), let us first prove that
$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
@@
where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ can be calculated by
@@
where the last equation holds due to (3) in Proposition 5.1. For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tend to identity in the sense of [37, Page 64, defintion], with a slight modification of the terminal time, such that
@@
where the convergence is understood as almost surely convergence. Therefore,
@@
where the last equation holds due to (3) in Proposition 5.1. Similar as the proof of (2c), we can conclude that
@@
$[X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{\mathbb{E}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
Now let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ being a sequence of partitions tend to identity in the sense of [37, Page 64, defintion] with a slight modification of the terminal time, such that
@@
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}),$
To repeat the same argument as before, we need to notice that $\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}$ and
[…]
where the first inequality is due to Cauchy inequality and Kunita-Watanabe Inequality and $C$ is a generic constant. Now repeating the arguments in (2c), we achieve

### v3

@@ 5. Proof of Results in Section 3
where the fourth equality holds by ${\mathcal{F}}_{t}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{t}$, the fifth equality follows from the same calculation as (5.5), and the last equality is by (2) in Proposition 5.1. Therefore,
$\overline{{\mathbb{P}}}(\widetilde{F}_{t}\cap\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t})=\overline{{\mathbb{P}}}(\widetilde{F}_{t}|\widetilde{\mathcal{G}}_{t})\overline{{\mathbb{P}}}(\widetilde{G}_{T}|\widetilde{\mathcal{G}}_{t}).$
5.2. Proof of Proposition 3.3
For (1), let us assume that ([51, Theorem 1, Chapter 4])
@@
where $M_{t}$ is an $\mathbb{F}$}-local martingale and $V_{t}$ is a finite variation process. To prove that $X$ is an $\overline{{\mathbb{F}}}$-semimartingale, it suffices to show that $M_{t}$ is an $\overline{{\mathbb{F}}}$-local martingale. Let $\{\tau_{k}\}_{k\in{\mathbb{N}}^{+}}$ be a sequence of $\mathbb{F}$}-stopping times such that $M^{\tau_{k}}:=(M_{t\wedge\tau_{k}})_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for all $k\in{\mathbb{N}}^{+}$, and $\tau_{k}\to\infty$ as $k\to\infty$. We claim that
@@
$\overline{{\mathcal{F}}}_{s}=\sigma\{A_{0}\times A_{1}\times A_{2}:A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}\},$
@@
for all $A_{0},A_{1},A_{2}\in{\mathcal{F}}_{s}$. Note that $M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}$, $\mathbbm{1}_{\Omega\times A_{1}\times\Omega}$ and $\mathbbm{1}_{\Omega\times\Omega\times A_{2}}$ are conditionally independent given ${\mathcal{G}}_{T}$ by (3) in Proposition 5.1, we have
@@
By (2) in Proposition 5.1 and ${\mathcal{F}}_{s}\perp\!\!\!\perp{\mathcal{G}}_{T}|{\mathcal{G}}_{s}$,
@@
$=\overline{{\mathbb{E}}}\Big{[}\overline{{\mathbb{E}}}\big{[}M_{t\wedge\tau_{k}}\mathbbm{1}_{A_{0}\times\Omega\times\Omega}\big{|}{\mathcal{G}}_{s}\big{]}\overline{{\mathbb{P}}}\big{(}A_{1}|{\mathcal{G}}_{s}\big{)}\overline{{\mathbb{P}}}\big{(}A_{2}|{\mathcal{G}}_{s}\big{)}\Big{]},$
where the second equality holds by the tower rule. Since $M^{\tau_{k}}$ is an $\mathbb{F}$}-martingale,
@@
Repeating the calculation as in (5.7), (5.8), and (5.9) yields
@@
To show that $X^{\prime}$ is an $\overline{{\mathbb{F}}}$-semimartingale, by the construction of the extended probability space, we can define $M^{\prime},V^{\prime}$, and $\tau_{k}^{\prime}$ to be respectively the conditionally independent copy of $M,V$, and $\tau_{k}$. Then it suffices to show that
@@
Note that
@@
where the first and last equalities mimic the argument in (5.7), (5.8) and (5.9), the second and the sixth equalities hold from (2) in Proposition 5.1, the third and the fifth equalities follow from the tower rule, and the fourth equality is the martingale property. This completes the proof of (1).
@@
For (2b), due to (2a) and the decomposability (see [51, Theorem 1, Chapter 4]), it suffices to show that if $X$ is an $\mathbb{F}$}-local martingale, then $X^{\mathcal{G}}$ is a local martingale with respect to $\overline{{\mathbb{F}}}$, $\mathbb{F}$} and $\mathbb{G}$}, which thanks to (1) can be further reduced to proving that $X^{\mathcal{G}}$ is a local martingale with respect to $\mathbb{F}$} and $\mathbb{G}$}. Let $\tau_{k}$ be the stopping time $\tau_{k}:=\inf\{t\in[0,T]:[X,X]_{t}>k\}$ and consider the truncated process $X^{\tau_{k}}:=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$. By Burkholder-Davis-Gundy inequality,
@@
for some constant $C$ and $(X^{\tau_{k}})^{*}_{t}=\sup_{0\leq s\leq t}|X_{s\wedge\tau_{k}}|.$ Let $\{\widetilde{\tau}_{\ell}\}_{\ell\in{\mathbb{N}}^{+}}$ be the sequence of the $\mathbb{F}$}-stopping times such that $\tau_{\ell}\to T$ and $X^{\tilde{\tau}_{\ell}}=\{X_{t\wedge\tilde{\tau}_{\ell}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$. Then $X^{\tilde{\tau}_{\ell}\wedge\tau_{k}}=\{X_{t\wedge\tilde{\tau}_{\ell}\wedge\tau_{k}}\}_{t\in[0,T]}$ is an $\mathbb{F}$}-martingale for any $\ell\in{\mathbb{N}}^{+}$. By Lemma 2.5, $|X_{t\wedge\tau_{k}\wedge\tilde{\tau}_{\ell}}|\leq(X^{\tau_{k}})^{*}_{t}$, and from (5.10),
@@
Combined with (5.10), we have that $X^{\tau_{k}}=\{X_{t\wedge\tau_{k}}\}_{t\in[0,T]}$ is indeed a square integrable $\mathbb{F}$}-martingale.
Now we claim that $(X^{\tau_{k}})^{\mathcal{G}}$ is a square integrable $\mathbb{G}$}-martingale. (Hence, by the H hypothesis and Remark 2.7 also an $\mathbb{F}$}-martingale). The integrability holds trivially since $X_{t}$ is integrable, and the martingale property holds since
@@
This completes the proof that $X^{\mathcal{G}}$ is an $\mathbb{G}$}-local martingale and (2b).
5.3. Proof of Theorem 3.5
For (1), let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions which tends to identity in the sense of [51, Page 64, definition] and satisfies
@@
where the convergence is understood as a.s. convergence. Then
@@
Recall that $X^{\prime}$ has the decomposition $X^{\prime}=M^{\prime}+V^{\prime}$, with $M^{\prime}$ a local martingale, $V^{\prime}$ an adapted finite variation process, and $\bigg{\|}\sqrt{[M^{\prime},M^{\prime}]_{T}}+\int_{0}^{T}\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\bigg{\|}_{L^{p}}<\infty.$ Therefore,
@@
For the first term in (5.11), applying Burkholder-Davis-Gundy inequality to get
@@
$\overline{{\mathbb{E}}}\Big{[}\Big{(}\sup_{0\leq s\leq t}(Z_{s})^{2}[M^{\prime},M^{\prime}]_{t}\Big{)}^{1/2}\Big{]}\leq\bigg{(}\overline{{\mathbb{E}}}\Big{[}\sup_{0\leq s\leq t}(Z_{s})^{q}\Big{]}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]^{p/2}_{t}\big{]}\bigg{)}^{1/2}\leq C\sqrt{\|Z\|^{q}_{{\mathcal{H}}^{q}}\overline{{\mathbb{E}}}\big{[}[M^{\prime},M^{\prime}]_{t}^{p/2}\big{]}}<\infty.$
Since $Z$ is càdlàg, $Z^{m}_{s}\to Z_{s-},a.s.$ and $|Z^{m}_{s}-Z_{s}|\leq\sup_{0\leq u\leq s}|Z_{u}|\in S_{q}$; and Lemma 2.5 leads to
@@
Note that $\int_{0}^{t}|Z_{s}^{m}-Z_{s-}|\big{|}{\mathrm{d}}V^{\prime}_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|\int_{0}^{t}|{\mathrm{d}}V_{t}|$, and similarly, we have
@@
$\overline{{\mathbb{P}}}\big{(}R_{m}-R>\epsilon\big{)}\leq\frac{\overline{{\mathbb{E}}}[(R_{m}-R)\mathbbm{1}_{A}]}{\epsilon}\to 0,$
@@
$\overline{{\mathbb{P}}}\big{(}R_{m}-R<\epsilon\big{)}\to 0.$
Therefore, $R_{m}\to R$ in probability. Meanwhile, $R_{m}\to\int_{0}^{t}Z_{s-}{\mathrm{d}}X^{\mathcal{G}}_{s}$ a.s., hence
@@
For (2), let us first prove that
$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{G}}\Big{]}.$
@@
where the last two terms are well defined because of (2b). The first term $X^{\mathcal{G}}_{t}Y^{\mathcal{G}}_{t}$ is
@@
where the last equation holds due to (3) in Proposition 5.1. For the second term $\int_{0}^{t}X^{\mathcal{G}}_{s-}{\mathrm{d}}Y^{\mathcal{G}}_{s}$, let $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ be a sequence of partitions tending to identity in the sense of [51, Page 64, definition] and satisfying
@@
where the convergence is understood as a.s. convergence. Therefore,
@@
where the last equation holds due to (3) in Proposition 5.1. Similar to the proof of (1), we conclude that
@@
$X^{\mathcal{G}},Y^{\mathcal{G}}]_{t}=\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},Y^{\prime}]_{t}\Big{|}{\mathcal{F}}\Big{]}.$
Meanwhile, we can further assume that the partition $\sigma_{m}:0=t_{0}^{m}\leq t_{1}^{m}\leq...\leq t_{k_{m}}^{m}=t$ satisfies
@@
$\int_{0}^{t}Z_{s-}{\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}=\lim_{m\to\infty}\sum_{k=0}^{k_{m}-1}Z_{t_{k}^{m}}([X^{\prime\prime},Y^{\prime}]_{t_{k+1}^{m}}-[X^{\prime\prime},Y^{\prime}]_{t_{k}^{m}}).$
Repeat the same argument as before and notice that
$\big{|}\int_{0}^{t}(Z_{s}^{m}-Z_{s-}){\mathrm{d}}[X^{\prime\prime},Y^{\prime}]_{s}\big{|}\leq 2\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{t}.$
Then we have
$\overline{{\mathbb{E}}}\bigg{[}\sup_{0\leq s\leq t}|Z_{s}|[X^{\prime\prime},Y^{\prime}]_{s}\bigg{]}$ $\leq\overline{{\mathbb{E}}}[(\sup_{0\leq s\leq t}|Z_{s}|)^{r}]^{1/r}\overline{{\mathbb{E}}}\Big{[}[X^{\prime\prime},X^{\prime\prime}]_{t}^{p}\Big{]}^{1/p}\overline{{\mathbb{E}}}\Big{[}[Y^{\prime},Y^{\prime}]_{t}^{q}\Big{]}^{1/q}$
$\leq C\|Z\|_{H_{r}}\|X\|_{H_{p}}\|Y\|_{H_{q}},$
where the first inequality is by Cauchy inequality and Kunita-Watanabe inequality with $C$ a generic constant. Now repeating the arguments in (1), we achieve